# Andreas Petrou-Zeniou

### v3

@@ Appendix A Alternative Approaches to Inference on Multiple Winners
In this section, I discuss three alternative approaches to the inference on multiple winners problem. Our first approach is novel, but closely related to the approach of Zrnic and Fithian (2024b). Our second approach is directly based on Zrnic and Fithian (2024b), while our final approach is based on Zrnic and Fithian (2024a). Our first approach uses a similar Bonferroni-type correction as the approach in section 5. However, this approach differs from our approach in section 5. Instead of estimating $\Delta$ from the first stage to model the errors on the winner in the second stage, we use $\Delta$ to estimate a set of likely winners, to which we restrict simultaneous inference in the second stage. This approach is thus very similar to that of Zrnic and Fithian (2024b), though it differs in a few subtle ways. We similarly demonstrate that the approach of Zrnic and Fithian (2024b) can be applied in our inference on multiple winners setting. Finally, we consider the approach of Zrnic and Fithian (2024a), who consider simultaneous testing of all means, but allocate the error budget according to the suboptimality (the observed gap of a given observation from the winner). Their approach is robust to arbitrary dependence, in the spirit of the Bonferroni and Holm (1979) procedures, and is thus conservative. Generally, all three of these approaches are underpowered compared to our two-step approach to inference.
A.1 A First Locally Simultaneous Approach to Inference
Let us consider the parametric setting from section 2. Let $\beta\in(0,\alpha)$, and let $\beta_{1}\in(0,\beta)$ with $\beta_{2}:=\beta-\beta_{1}$. For arbitrary $\mu$ and $\Sigma$, let us pick some $J_{c}\subseteq J$ be the smallest set such that: [^7] If multiple equal cardinality $J_{c}$ satisfy 22, we choose the $J_{c}$ minimizing the probability on the left hand side of 22.
@@ A.1 A First Locally Simultaneous Approach to Inference
Similarly, we seek to construct a prediction set $\hat{J}_{c}$ for $J_{c}$ such that $P_{\mu,\Sigma}\left(J_{c}\subseteq\hat{J}_{c}\right)\geq 1-\beta_{2}$. Finally, recall that in section 4, we defined $c_{1-\alpha}(J_{c})$ to be the $1-\alpha$-quantile of 7.
It remains to show how to construct $\hat{J}_{c}$. By 8, it follows that:
$p_{j}:=P_{\mu,\Sigma}\left(j\in\hat{J}_{R}\right)=P_{\mu,\Sigma}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+\Delta_{j^{\prime},j}\right)\in R\right)$ (23)
Since the $\Delta_{j}$ are unknown, we cannot compute the above probabilities. Instead, we take the familiar approach of computing upper and lower confidence bounds for the $\Delta_{j}$, which we denote by $U$ and $L$ respectively, at level $1-\beta_{2}$. We constructed such $U$ and $L$ in equation 10.
It follows that $P_{\mu,\Sigma}(L\leq\Delta\leq U)\geq 1-\beta_{2}$. We can now observe that this probability no longer depends on $\mu$. Let $p_{j}(L,U)$ be the right hand side of 23 with $L$ and $U$ in place of $\Delta$, which we take as follows:
$p_{j}(L,U)=P_{\mu,\Sigma}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+U_{j^{\prime},j}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+L_{j^{\prime},j}\right)\right]\cap R\neq\emptyset\right)$ (24)
We have that $P_{\mu,\Sigma}\left(p(L,U)\leq p\right)\geq 1-\beta_{2}$, where inequality is interpreted elementwise. We take $\hat{J}_{c}$ to be the set that minimizes $\sum_{j\in\hat{J}_{c}}p_{j}(L,U)$ subject to the requirement that this sum exceed $1-\beta_{1}$. If no such set exists, we take $\hat{J}_{c}=J$.
@@
$CS_{1-\alpha;\beta}^{LS1}:=\times_{\jmath\in\hat{J}_{R}}\left[Y_{\jmath}-c_{1-(\alpha-\beta)}\left(\hat{J}_{c}\right)\sqrt{\Sigma_{Y,\jmath\jmath}},\ Y_{\jmath}+c_{1-(\alpha-\beta)}\left(\hat{J}\right)\sqrt{\Sigma_{Y,\jmath\jmath}}\right]$
@@
$CS_{1-\alpha;\beta}^{LS1}$ is a valid confidence set at the $1-\alpha$-level, such that:
$P_{\mu,\Sigma}\left(\left(\mu_{Y,\jmath}\right)_{\jmath\in\hat{J}_{R}}\in CS_{1-\alpha;\beta}^{LS1}\right)\geq 1-\alpha$ (25)
@@
A.2 A Second Locally Simultaneous Approach to Inference
We now adapt the locally simultaneous approach of Zrnic and Fithian (2024b) to the inference on multiple winners problem. As before, Zrnic and Fithian (2024b) construct some set of likely selections $\hat{J}_{c}$ and conduct simultaneous inference restricted to this $\hat{J}_{c}$. In general, Zrnic and Fithian (2024b) construct this $\hat{J}_{c}$ by first constructing a $1-\beta$ confidence region for the data generating process $P$. In our case, we can equivalently construct a $1-\beta$ confidence region for $\mu$, or some other carefully-chosen parameter of interest. Subsequently, for each $P$ in the confidence region described above, Zrnic and Fithian (2024b) derive a $1-\beta$-forecast set for the observations generated by this $P$, and the selections these observations imply. Taking the union of these forecast sets over all $P$ in the aforementioned confidence set yields a forecast set for the selection which is valid at level $1-\beta$. This forecast set will provide the $\hat{J}_{c}$ described above.
For the inference on multiple winners problem, we take:
$\hat{J}_{c}:=\left\{j;X_{j}\not\in\bigcup_{\jmath\in\hat{J}_{R}}\left[X_{\jmath}-2d_{1-\beta/2}(\Sigma),\ X_{\jmath}+2d_{1-\beta/2}(\Sigma)\right]\right\}$ (26)
Finally, we may define the following confidence set:
$CS_{1-\alpha;\beta}^{LS2}:=\times_{\jmath\in\hat{J}_{R}}\left[Y_{\jmath}-c_{1-(\alpha-\beta)}\left(\hat{J}_{c}\right)\sqrt{\Sigma_{Y,\jmath\jmath}},\ Y_{\jmath}+c_{1-(\alpha-\beta)}\left(\hat{J}\right)\sqrt{\Sigma_{Y,\jmath\jmath}}\right]$
As before, we have the following proposition:
Proposition A.2.
$CS_{1-\alpha}^{LS1}$ is a valid confidence set at the $1-\alpha$-level, such that:
$P_{\mu,\Sigma}\left(\left(\mu_{Y,\jmath}\right)_{\jmath\in\hat{J}_{R}}\in CS_{1-\alpha}^{LS2}\right)\geq 1-\alpha$ (27)
for all $\mu$ and $\Sigma$. Again, marginal validity clearly follows.
A proof of this proposition is provided in appendix C.
@@ A.3 A Zoom Test for Inference
In this section, we provide an approach to inference on multiple winners based on the zoom test of Zrnic and Fithian (2024a). Their approach suggests allocating the error budget to near-winners by inverting a test which Zrnic and Fithian (2024b) call the zoom test. The zoom test is based on an acceptance region which is increasing in the population suboptimality - the difference between a candidate’s mean and the population winner’s mean. While it is unclear how to generalize their approach to the exact inference on multiple winners problem we discuss in this paper, they provide guidance
@@
Clearly, under the point hypothesis $H_{0}(\mu_{Y},\mu_{X}):\mathbb{E}(Y)=\mu_{Y},\ \mathbb{E}(X)=\mu_{X}$, the probability that $(Y_{j})_{j\in J}$ lies in $A_{\alpha}$ exceeds $1-\alpha$. We define the following, joint confidence for $\left(\mu_{Y,\jmath}\right)_{\jmath\in\hat{J}_{R}}$:
@@
For any $\jmath\in\hat{J}_{R}$, we define the marginal confidence set:
$CS_{1-\alpha,\jmath}^{zoom}:=\left\{\mu_{Y,\jmath};\exists\ \tilde{\mu}_{Y},\ \tilde{\mu}_{X}\in CS_{1-\alpha}\quad\text{s.t.}\quad\tilde{\mu}_{Y,\jmath}=\mu_{Y,\jmath}\right\}$
@@
$CS_{1-\alpha}^{zoom}:=\times_{\jmath\in\hat{J}_{R}}CS_{1-\alpha,\jmath}^{zoom}$ (30)
Of course, it is not clear how to invert the test based on the acceptance region $A_{\alpha}$. Zrnic and Fithian (2024a) provide computationally efficient procedures to invert the zoom test under certain special cases, namely when $R$ is of the form $\{1,\ldots,\tau\}$ for some $\tau\leq p$. Zrnic and Fithian (2024b) provide a step-wise implementation of their methods which we implement in all simulations included in this paper.
@@ A.4 Simulations Comparing Different Approaches to Inference
We now provide results from a small simulation study comparing our two-step approach to inference with the approaches above. We provide results from a more extensive simulation study in section 13. In the simulations below, we consider $\Sigma=I_{2}$, $\mu=\begin{pmatrix}\mu_{1}&0\end{pmatrix}^{\prime}$, and vary $\mu_{1}$. Our two-step approach to inference substantially outperforms the approaches of Zrnic and Fithian (2024b) and Zrnic and Fithian (2024a) for intermediate values of $\mu_{1}$. In particular, our approach reduces over-coverage error by as much as 68.6%. However, for extreme values of $\mu_{1}$, where there winner is clear, the approaches of Zrnic and Fithian (2024b) and Zrnic and Fithian (2024a) slightly outperform our two-step approach, reducing over-coverage by up to 43.4%, although over-coverage for all methods is low in such instances. Our simulation results are plotted below:
@@
A more comprehensive simulation study is provided in appendix E.
@@ Appendix B Uniform Asymptotic Validity and Proofs
In this section, we formalize the assumptions stated in section 5.2 and introduce several lemmas supporting the proof of proposition B.1. In particular, the lemmas in this section closely resemble proposition 11 from Andrews et al. (2023), in that we derive uniform asymptotic validity results on conservative, rectangular confidence intervals.
B.1 Assumptions
In this section, we provide a list of formally-stated assumptions necessary to prove propositions B.1 and 5.3. First, we provide the following uniform integrability assumption, which is equivalent to uniform convergence in distribution, per lemma B.1.
@@ B.1 Assumptions
$\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left(\frac{\left|\tilde{W}_{1,j}-\mu_{W,j}(P)\right|^{2}}{\Sigma_{W,jj}(P)}\mathbb{1}\left(\frac{\left|\tilde{W}_{1,j}-\mu_{W,j}(P)\right|}{\sqrt{\Sigma_{W,jj}(P)}}>K\right)\right)<\varepsilon$
Moreover, we impose the following uniform consistency condition on our estimator $\hat{\Sigma}^{n}$ for $\Sigma(P)$.
Assumption B.2.
We assume that $\hat{\Sigma}^{n}$ is a uniformly, asymptotic consistent estimator for $\Sigma(P)$, such that, for any $\varepsilon>0$:
$\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left\|\hat{\Sigma}^{n}-\Sigma(P)\right\|>\varepsilon\right)=0$
where our choice of matrix norm is the max norm.
@@
Under assumptions B.1-B.2, we have uniform convergence in distribution such that there exist functions $\mu_{W,n}(P):=\mu_{W}(P)$ satisfying:
@@
Our second lemma concerns the asymptotic properties of $L^{n}$, and namely, whether $L^{n}$ satisfies a uniform, asymptotic validity condition as a lower bound of $\Delta^{n}_{j^{\prime},j}\equiv\Delta^{n}_{j^{\prime},j}(P):=\mu_{X,j^{\prime}}(P)-\mu_{X,j}(P)$. We naturally define $L^{n}$ and $U^{n}$ as follows:
$L_{j^{\prime},j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}-d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\frac{\hat{\text{var}}^{n}_{j^{\prime}j}}{n}}$
$U_{j^{\prime},j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}+d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\frac{\hat{\text{var}}^{n}_{j^{\prime}j}}{n}}$
where $\hat{\text{var}}^{n}_{j^{\prime},j}=\hat{\Sigma}^{n}_{j^{\prime}j^{\prime}}+\hat{\Sigma}^{n}_{jj}-2\hat{\Sigma}^{n}_{j^{\prime}j}$, with an analogous defintion for $\text{var}_{j^{\prime},j}(P)$.
@@
Given assumption B.1, it follows that $L^{n}$ is a uniformly, asymptotically valid lower bound for $\Delta^{n}(P)$. In particular:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(L^{n}\leq\Delta^{n}(P)\leq U^{n}\right)\geq 1-\beta$

### v4

@@ Appendix A Alternative Approaches to Inference on Multiple Winners
In this section, we discuss three alternative approaches to the inference on multiple winners problem. We first show that our two-step approach to inference can be intersected with a projection confidence set in the spirit of theorem 2 Zrnic and Fithian (2024b). Our second approach to inference is novel, but closely related to the approach of Zrnic and Fithian (2024b). Our final approach is based on Zrnic and Fithian (2024a). Our first approach uses a similar Bonferroni-type correction as the approach in section 3. However, this approach differs from our approach in section 3. Instead of estimating $\Delta$ from the first stage to model the errors on the winner in the second stage, we use $\Delta$ to estimate a set of likely winners, to which we restrict simultaneous inference in the second stage. This approach can be thought of as a version of the locally simultaneous inference of Zrnic and Fithian (2024b). We similarly demonstrate that the approach of Zrnic and Fithian (2024b) can be applied in our inference on multiple winners setting. Finally, we consider the approach of Zrnic and Fithian (2024a), who consider simultaneous testing of all means, but allocate the error budget according to the suboptimality (the observed gap of a given observation from the winner). Generally, all three of these approaches are underpowered compared to our two-step approach to inference.
A.1 A weakly improved two-step approach
We provide the following non-inferiority result in the spirit of theorem 2 of Zrnic and Fithian (2024b). To do so, we will introduce some new notation. For any subset $J_{c}$ of $J$, we define $\bar{c}_{1-\alpha}(J_{c})$ to be the $(1-\alpha)$-quantile of:
$CS_{1-\alpha;\beta}^{TS2}:=\times_{\hat{\jmath}\in\hat{J}_{R}}\left[Y_{j}-\left(\rho_{1-\alpha+\beta}(L,U)\wedge\bar{c}_{1-\alpha}\right)\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}},Y_{j}+\left(\rho_{1-\alpha+\beta}(L,U)\wedge\bar{c}_{1-\alpha}\right)\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}\right$]
Suppose that $X=Y$. Then $CS_{1-\alpha;\beta}^{TS2}\subseteq CS_{1-\alpha}^{P}$.
Proposition A.1.
Suppose $\beta$ is sufficiently small such that $2\bar{c}_{1-\alpha}\leq d_{1-\beta}(\Sigma)$. Then $CS_{1-\alpha;\beta}^{TS2}$ is a valid confidence set at the $1-\alpha$-level, such that:
$P_{\mu,\Sigma}\left(\left(\mu_{Y,\hat{\jmath}}\right)_{\hat{\jmath}\in\hat{J}_{R}}\in CS_{1-\alpha;\beta}^{TS2}\right)\geq 1-\alpha$ (26)
for all $\mu$ and $\Sigma$.
We prove this result in appendix C.
A.2 A Two-Step, Locally Simultaneous Approach to Inference
Let us consider the parametric setting from section 2. Let $\beta\in(0,\alpha)$, and let $\beta_{1}\in(0,\beta)$ with $\beta_{2}:=\beta-\beta_{1}$. For arbitrary $\mu$ and $\Sigma$, let us pick some $J_{c}\subseteq J$ to be the smallest set such that: [^8] If multiple $J_{c}$ of equal cardinality satisfy (27), we choose the $J_{c}$ minimizing the probability on the left hand side of (27).
@@ A.2 A Two-Step, Locally Simultaneous Approach to Inference
Similarly, we seek to construct a confidence set $\hat{J}_{c}$ for $J_{c}$ such that $P_{\mu,\Sigma}\left(J_{c}\subseteq\hat{J}_{c}\right)\geq 1-\beta_{2}$. Finally, recall that in section 4, we defined $c_{1-\alpha}(J_{c})$ to be the $1-\alpha$-quantile of (19).
It remains to show how to construct $\hat{J}_{c}$. By (4), it follows that:
$p_{j}:=P_{\mu,\Sigma}\left(j\in\hat{J}_{R}\right)=P_{\mu,\Sigma}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+\Delta_{j^{\prime}j}\right)\in R\right)$ (28)
Since the $\Delta_{j}$ are unknown, we cannot compute the above probabilities. Instead, we take the familiar approach of computing upper and lower confidence bounds for the $\Delta_{j}$, which we denote by $U$ and $L$ respectively, at level $1-\beta_{2}$. We constructed such $U$ and $L$ in equation (6).
It follows that $P_{\mu,\Sigma}(L\leq\Delta\leq U)\geq 1-\beta_{2}$. We define $p_{j}(l,u)$ as follows:
$p_{j}(l,u)=P_{\mu,\Sigma}\left(\left[\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+u_{j^{\prime}j}\right),\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{X,j}\geq\xi_{X,j^{\prime}}+l_{j^{\prime},j}\right)\right]\cap R\neq\emptyset\right)$
We have that $P_{\mu,\Sigma}\left(p(L,U)\leq p\right)\geq 1-\beta_{2}$, where the inequality is interpreted elementwise. We take $\hat{J}_{c}$ to be the set that minimizes $\sum_{j\in\hat{J}_{c}}p_{j}(L,U)$ subject to the requirement that this sum exceed $1-\beta_{1}$. If no such set exists, we take $\hat{J}_{c}=J$.
@@
$CS_{1-\alpha;\beta}^{LSTS}:=\times_{\hat{\jmath}\in\hat{J}_{R}}\left[Y_{\hat{\jmath}}-c_{1-(\alpha-\beta)}\left(\hat{J}_{c}\right)\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}},\ Y_{\hat{\jmath}}+c_{1-(\alpha-\beta)}\left(\hat{J}\right)\sqrt{\Sigma_{Y,\hat{\jmath}\hat{\jmath}}}\right$]
@@
$CS_{1-\alpha;\beta}^{LSTS}$ is a valid confidence set at the $1-\alpha$-level, such that:
$P_{\mu,\Sigma}\left(\left(\mu_{Y,\hat{\jmath}}\right)_{\hat{\jmath}\in\hat{J}_{R}}\in CS_{1-\alpha;\beta}^{LSTS}\right)\geq 1-\alpha$ (29)
@@ A.3 A Zoom Test for Inference
In this section, we provide an approach to inference on multiple winners based on the zoom test of Zrnic and Fithian (2024a). Their approach suggests allocating the error budget to near-winners by inverting the zoom test. The zoom test is based on an acceptance region which is increasing in the population suboptimality - the difference between a candidate’s mean and the population winner’s mean. While it is unclear how to precisely generalize their approach to the exact inference on multiple winners problem we discuss in this paper, Zrnic and Fithian (2024b) provide guidance on the case of selecting the top-$\tau$ winners.
@@
Clearly, under the point hypothesis $H_{0}(\mu_{Y},\mu_{X}):\mathbb{E}(Y)=\mu_{Y},\ \mathbb{E}(X)=\mu_{X}$, the probability that $(Y_{j})_{j\in J}$ lies in $A_{\alpha}$ exceeds $1-\alpha$. We define the following, joint confidence for $\left(\mu_{Y,\hat{\jmath}}\right)_{\hat{\jmath}\in\hat{J}_{R}}$:
@@
For any $\hat{\jmath}\in\hat{J}_{R}$, we define the marginal confidence set:
$CS_{1-\alpha,\hat{\jmath}}^{zoom}:=\left\{\mu_{Y,\hat{\jmath}};\exists\ \tilde{\mu}_{Y},\ \tilde{\mu}_{X}\in CS_{1-\alpha}\quad\text{s.t.}\quad\tilde{\mu}_{Y,\hat{\jmath}}=\mu_{Y,\hat{\jmath}}\right\}$
@@
$CS_{1-\alpha}^{zoom}:=\times_{\hat{\jmath}\in\hat{J}_{R}}CS_{1-\alpha,\hat{\jmath}}^{zoom}$ (32)
Of course, it is not clear how to invert the test based on the acceptance region $A_{\alpha}$ for completely general $R$. Zrnic and Fithian (2024a) provide a parsimonious characterization of the confidence set based on inversion of the zoom test under certain cases, namely when $R$ is of the form $\{1,\ldots,\tau\}$ for some $\tau\leq p$. Computationally, Zrnic and Fithian (2024b) provide a step-wise implementation of their methods which we implement in all simulations included in this paper.
@@ A.4 Simulations Comparing Different Approaches to Inference
We now provide results from a small simulation study comparing our two-step approach to inference with the approaches above. We provide results from a more extensive simulation study in section 11. In the simulations below, we consider $\Sigma=I_{2}$, $\mu=\begin{pmatrix}\mu_{1}&0\end{pmatrix}^{\prime}$, and vary $\mu_{1}$. Our two-step approach to inference substantially outperforms the approaches of Zrnic and Fithian (2024b) and Zrnic and Fithian (2024a) for intermediate values of $\mu_{1}$. In particular, our approach reduces over-coverage error by as much as 68.6%. However, for extreme values of $\mu_{1}$, where the winner is clear, the approaches of Zrnic and Fithian (2024b) and Zrnic and Fithian (2024a) slightly outperform our two-step approach, reducing over-coverage by up to 43.4%, although over-coverage for all methods is low in such instances. Our simulation results are plotted below:
@@
A more comprehensive simulation study is provided in appendix D.
@@ Appendix B Uniform Asymptotic Validity and Proofs
In this section, we formalize the assumptions stated in section 3.2 and introduce several lemmas supporting the proof of proposition 3.2. First, we provide the following uniform integrability assumption, which is equivalent to uniform convergence in distribution, per lemma B.1.
@@
$\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left(\frac{\left|\tilde{W}_{1,j}-\mu_{W,j}(P)\right|^{2}}{\Sigma_{W,jj}(P)}\mathbb{1}\left(\frac{\left|\tilde{W}_{1,j}-\mu_{W,j}(P)\right|}{\sqrt{\Sigma_{W,jj}(P)}}>K\right)\right)<\varepsilon$ (33)
@@
Under assumption B.1, we have uniform convergence in distribution such that there exist functions $\mu_{W,n}(P):=\mu_{W}(P)$ satisfying:
@@
Our second lemma provides a useful consequence of assumption B.1.
@@
Under assumption B.1, it holds that for $j,j^{\prime}=1,\ldots,2p$ such that $j\neq j^{\prime}$ and such that $\text{var}_{jj^{\prime}}(P)\neq 0$, and for any $\varepsilon$, there exists $K$ sufficiently large such that:
$\sup_{P\in\mathcal{P}}\mathbb{E}_{P}\left(\frac{\left|(\tilde{W}_{1,j}-\mu_{W,j}(P))-(\tilde{W}_{1,j^{\prime}}-\mu_{W,j^{\prime}}(P))\right|^{2}}{\text{var}_{jj^{\prime}}(P)}\mathbb{1}\left(\frac{\left|(\tilde{W}_{1,j}-\mu_{W,j}(P))-(\tilde{W}_{1,j^{\prime}}-\mu_{W,j^{\prime}}(P))\right|}{\sqrt{\text{var}_{jj^{\prime}}(P)}}>K\right)\right)<\varepsilon$